# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v1

@@ III-A Essentials on RES distributions
where the random vector $\mathbf{u}\sim\mathcal{U}(S_{\mathbb{R}}^{m-1})$ is uniformly distributed on the unit sphere $S_{\mathbb{R}}^{m-1}\triangleq\{\mathbf{u}\in\mathbb{R}^{m}|||\mathbf{u}||=1\}$ and consequently satisfies $E\{\mathbf{u}\}=\mathbf{0}$ and $E\{\mathbf{u}\mathbf{u}^{T}\}=m^{-1}\mathbf{I}_{m}$. The positive random variable $\mathcal{Q}$, called 2nd-order modular variate, is such that (s.t.)
@@
Note that, from the stochastic representation (31), the properties of $\mathbf{u}\sim S_{\mathbb{R}}^{m-1}$ and the fact that $\mathcal{Q}$ is independent of $\mathbf{u}$, the constraint $E\{\mathcal{Q}\}=m$ implies that
@@
where $\mathcal{S}_{m,S}^{\mathbb{R}}$ is a (non-linear) differentiable manifold on dimension $m(m+1)/2-1$ such that:
@@
We note in passing that the most popular choices for the scale function $S(\cdot)$ are $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$ and $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$. It is worth emphasizing that, under A1, A2 and A3, and as a direct consequence of the implicit function theorem, the first top-left entry of $\mathbf{V}_{S}$, i.e. $[\mathbf{V}_{S}]_{11}$, can (locally at $\mathbf{V}_{S}$) be expressed as function of the other entries. For example, if we choose $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, then $[\mathbf{V}_{S,0}]_{11}$ is trivially given by $[\mathbf{V}_{S,0}]_{11}=1$. For $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$, we have that $[\mathbf{V}_{S}]_{11}=m-\sum\nolimits_{i=2}^{m}[\mathbf{V}_{S}]_{ii}$. Lastly, for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$ it can be easily shown that, using the Laplace’s expansion of the determinant along the first row of $\mathbf{V}_{S}$, $[\mathbf{V}_{S}]_{11}$ can be recovered as $[\mathbf{V}_{S}]_{11}=\frac{1}{C_{11}}\left(1-\sum\nolimits_{i=2}^{m}[\mathbf{V}_{S}]_{1i}C_{1i}\right)$ where $C_{ij}$ indicates the cofactor of $[\mathbf{V}_{S}]_{ij}$.
@@ III-B The model $\mathcal{P}_{\bm{\eta},g}$ and its parametric and semiparametric submodels
as the finite dimensional parameter vector of interest, we can define the following parametric submodel of $\mathcal{P}_{\bm{\eta},g}$:
@@ III-C The score vectors and the related FIM in $\mathcal{P}_{\bm{\eta},g}$
where $s_{0}=S(\bm{\Sigma}_{0})$ and $\mathbf{V}_{S,0}=\bm{\Sigma}_{0}/s_{0}$. Furthermore, the 2nd diagonal component in (51), as shown in [10, Sect. 4] and, by using our notation in the Appendix 156, of this work is given by
@@
where $S([\mathbf{V}_{S}]_{11},\underline{\mathrm{vecs}}(\mathbf{V}_{S}))$ denotes the scale function applied to $\mathbf{V}_{S}$ reconstructed from $[\mathbf{V}_{S}]_{11}$ and $\underline{\mathrm{vecs}}(\mathbf{V}_{S})$. For further reference, let us recast $\mathbf{J}[\mathbf{w}](\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0})$ as:
@@
evaluated at $\mathbf{V}_{S,0}$. As an example, we show in Appendix 162 that the term $\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}$ can be explicitly obtained for the three above-mentioned scale functions as:
@@
  • $-\frac{\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{(\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m})_{1}}$ for $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$,
  • $-\frac{\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{(\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m})_{1}}$ for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$.
@@
Let us now introduce the score vector for $\bm{\nu}_{0}\in\Omega$ in the model $\mathcal{P}_{\bm{\nu},g}$ in (39) as $\mathbf{s}_{\bm{\nu}_{0}}=(\mathbf{s}_{\bm{\mu}_{0}}^{T},\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{T})^{T}$. By applying the rules of the differential matrix calculus detailed in [23, Ch. 8], the Kronecker product and the vec operator [23, Ch. 3], and the stochastic representation theorem (31), we get the following expressions of the scores (see [22, eqs. (34) and (37)]):
@@
where, following the notation adopted in [8], we defined:  [^2] To avoid confusion, it is important to note that the function $\varphi_{0}(t)$ is linked to the equivalent function $\psi_{0}(t)$ used in [22, 24] by the constant $-2$: $\varphi_{0}(t)=-2\psi_{0}(t)$.
@@
[…]
@@
In this section, we will specialise Lemmas II.1 and II.2, introduced in Sect. II, to the study of the elliptical distributions. The semiparametric model from which we start is the “canonical” one, i.e. $\mathcal{P}_{\bm{\eta},g}$ given in (40), corresponding to (c.t.) $\mathcal{P}$ in (5), along with its parametric submodels $\mathcal{P}_{\bm{\gamma}}$ in (44) (c.t. $\mathcal{P}_{1}$ in (6)), $\mathcal{P}_{s}$ in (45) (c.t. $\mathcal{P}_{2}$ in (7)), $\mathcal{P}_{\bm{\gamma},s}$ in (46) (c.t. $\mathcal{P}_{1,2}$ in (8)) and the non-parametric model $\mathcal{P}_{g}$ in (47) (c.t. $\mathcal{P}_{3}$ in (9)). Moreover, we will investigate the link of the parametric model $\mathcal{P}_{\gamma,s,\bm{\zeta}}$, introduced in (48), with the above-mentioned models. The key result that allows us to study the impact of finite and infinite dimensional nuisance parameters on the semiparametric efficiency in $\mathcal{P}_{\bm{\eta},g}$ for the parameter vector of interest $\bm{\gamma}_{0}$ is Lemma II.2. To specialize this lemma to our context, we need to draw the following connections:
  • the finite-dimensional vector of the parameters of interest is $\bm{\gamma}_{0}=(\bm{\mu}_{0}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})^{T})^{T}$ and the related score
@@
  • The infinite-dimensional nuisance tangent $\mathcal{T}_{g_{0}}$ (c.t. $\mathcal{T}_{3}$) is given in Appendix .7 as:
@@
$\bar{\mathbf{t}}_{\bm{\gamma}_{0}}=\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})\perp\mathcal{T}_{g_{0}},$ (77)
@@
It now follows from equations (80) and (81) that condition (78) is satisfied. then, the Condition (78) is satisfied.
@@
or equivalently $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$ iff $\mathrm{vec}(\mathbf{V}_{S,0}^{-1})\in\text{Ker}\;\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})=\{a\cdot\mathrm{vec}(\mathbf{D}_{S}^{\mathbf{V}_{S,0}}),a\in\mathbb{R}\}$. As already noted in [10, Th. 3.1], this is a direct consequence of the Property P4 of the matrix $\mathbf{M}_{S}^{\mathbf{V}_{S,0}}$ (see appendix .6). In fact, for $S\equiv S_{d}$, we have that $\mathbf{D}_{S_{d}}^{\bm{\Sigma}_{0}}=m^{-1}|\bm{\Sigma}_{0}|^{1/m}\bm{\Sigma}_{0}^{-1}=m^{-1}\mathbf{V}_{S,0}^{-1}$. Consequently, we immediately have that $\mathrm{vec}(\mathbf{V}_{S_{d},0}^{-1})\in\text{Ker}\;\mathbf{M}_{S_{d}}^{\mathbf{V}_{S,0}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$ and then $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$. ∎
@@
$\bar{\mathbf{t}}_{\bm{\gamma}_{0}}=\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})\perp\mathcal{T}_{\bm{\zeta}_{0}},$ (86)
@@
V Parametric and semiparametric information bounds in $\mathcal{P}_{\bm{\eta},g}$ and $\mathcal{P}_{\bm{\nu},g}$
@@
However, as said before, the canonical parameter of interest $\bm{\gamma}_{0}$ contains the shape matrix $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$ that belongs to the differentiable but non-linear manifold $\mathcal{S}_{m,S}^{\mathbb{R}}$ in (38) and this causes considerable difficulties when conducting inference on $\mathbf{V}_{S,0}$. For this reason, it is often preferable to make a “constrained” inference in the model $\mathcal{P}_{\bm{\nu},g}$ in (39) where the parameter $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma}_{0})^{T})^{T}$ contains the scatter matrix $\bm{\Sigma}_{0}$ belonging to $\mathcal{S}_{m}^{\mathbb{R}}$ that is indeed a linear subspace of $\mathbb{R}^{m\times m}$. Clearly, performing inference on $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$ in $\mathcal{P}_{\bm{\eta},g}$ in the presence of the nuisance parameters $s_{0}$ and $g_{0}$ is equivalent (at least locally at $\mathbf{V}_{S,0}$, thanks to the implicit function theorem) performing inference on ${\bm{\Sigma}}_{0}\in\mathcal{S}_{m}^{\mathbb{R}}$ in $\mathcal{P}_{\bm{\nu},g}$ under the constraint $S(\bm{\Sigma}_{0})=s_{0}$ satisfying the assumptions A1, A2 and A3 in Sect. III. This highlights the usefulness of the two constrained information/CR bounds:
  • Constrained parametric CRB (CCRB) for the parameters $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma}_{0})^{T})^{T}$ under the constraint $S(\bm{\Sigma}_{0})=s_{0}$ in the parametric submodel $\mathcal{P}_{\bm{\nu},g_{0}}$. This bound is well-defined in the standard parametric theory as shown in [26].
  • Constrained Semiparametric CRB (CSCRB) for the parameters $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma}_{0})^{T})^{T}$ under the constraint $S(\bm{\Sigma}_{0})=s_{0}$ in the semiparametric submodel $\mathcal{P}_{\bm{\nu},g}$ in the presence of the infinite-dimensional nuisance $g_{0}\in\bar{\mathcal{G}}$. This bound is well-defined by a constrained generalization of the Hájek-Le Cam convolution theorem as shown in [27].
The following three subsections are organized as follows: in subsect. V-A, an explicit expression for the parametric CRB of $\bm{\gamma}_{0}$ in $\mathcal{P}_{\bm{\gamma},s}$ in the presence of the nuisance parameter $s_{0}$ is provided, while in subsect. V-B the constrained parametric CRB for $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma}_{0})^{T})^{T}$ under the constraint $S(\bm{\Sigma}_{0})=s_{0}$ is given. Finally, the subsect. V-C proposes the counterpart of the Propositions IV.1 and IV.2 in terms of parametric and semiparametric information bounds.
@@
To calculate (in a closed-form not containing matrix inverses) the $\mathrm{CRB}(\bm{\gamma}_{0}|s_{0})$ in (91), we will exploit the expression of the CRB on $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma})_{0}^{T})^{T}$ deduced from the inverse of the FIM $\mathbf{I}(\bm{\nu}_{0})$. To this end, let us consider the inverse, say $\mathbf{w}^{-1}$, of the diffeomorphism $\mathbf{w}$ given in (50). It can be explicitly obtained as:
$\begin{split}\mathbf{w}^{-1}:\;&\Omega\rightarrow\Phi\\ &\bm{\nu}=(\bm{\mu}^{T},\mathrm{vecs}(\bm{\Sigma})^{T})^{T}\mapsto\mathbf{w}^{-1}(\bm{\nu})=(\bm{\mu}^{T},S^{-1}(\bm{\Sigma})\underline{\mathrm{vecs}}(\bm{\Sigma})^{T},S(\bm{\Sigma}))^{T},\end{split}$ (94)
@@
The following proposition is proved in Appendix 199:
@@
$\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})\triangleq\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}=\alpha^{-1}(g_{0})\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\mathbf{P}_{S}^{T}(\mathbf{V}_{S,0})\mathbf{D}_{m}^{\#T}\underline{\mathbf{I}}_{m}^{T},$ (98)
@@
  • $\mathbf{I}_{m^{2}}-{\rm vec}(\mathbf{V}_{S,0}){\bf e}^{T}_{1}$ for $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$,
  • $\mathbf{I}_{m^{2}}-\frac{1}{m}{\rm vec}(\mathbf{V}_{S,0}){\rm vec}(\mathbf{I}_{m^{2}})^{T}$ for $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$,
  • $\mathbf{I}_{m^{2}}-\frac{1}{m}{\rm vec}(\mathbf{V}_{S,0}){\rm vec}(\mathbf{V}_{S,0}^{-1})^{T}$ for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$.

### v2

@@ III-A Essentials on RES distributions
[…]
@@
Note that, from the stochastic representation (31), the properties of $\mathbf{u}\sim\mathcal{U}(S_{\mathbb{R}}^{m-1})$ and the fact that $\mathcal{Q}$ is independent of $\mathbf{u}$, the constraint $E\{\mathcal{Q}\}=m$ implies that
@@
where $\mathcal{S}_{m,S}^{\mathbb{R}}$ is a (non-linear) differentiable manifold on dimension m(m+1)/21m(m+1)/2-1 such that:
@@
We note in passing that the most popular choices for the scale function $S(\cdot)$ are $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$ and $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$. It is worth emphasizing that, under A1, A2 and A3, and as a direct consequence of the implicit function theorem, the first top-left entry of $\mathbf{V}_{S}$, i.e. $\mathbf{V}_{S}]_{11}$, can (locally at $\mathbf{V}_{S}$) be expressed as function of the other entries. For example, if we choose $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, then $\mathbf{V}_{S,0}]_{11}$ is trivially given by $\mathbf{V}_{S}]_{11}=1$. For $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$, we have that $\mathbf{V}_{S}]_{11}=m-\sum\nolimits_{i=2}^{m}[\mathbf{V}_{S}]_{ii}$. Lastly, for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$ it can be easily shown that, using the Laplace’s expansion of the determinant along the first row of $\mathbf{V}_{S}$, $\mathbf{V}_{S}]_{11}$ can be recovered as $\mathbf{V}_{S}]_{11}=\frac{1}{C_{11}}\left(1-\sum\nolimits_{i=2}^{m}(-1)^{1+i}[\mathbf{V}_{S}]_{1i}C_{1i}\right)$ where $C_{ij}$ indicates the cofactor of $\mathbf{V}_{S}]_{ij}$.
@@ III-B The model $\mathcal{P}_{\bm{\eta},g}$ and its parametric and semiparametric submodels
the finite dimensional parameter vector of interest, we can define the following parametric submodel of $\mathcal{P}_{\bm{\eta},g}$:
@@ III-C The score vectors and the related FIM in $\mathcal{P}_{\bm{\eta},g}$
where $s_{0}=S(\bm{\Sigma}_{0})$ and $\mathbf{V}_{S,0}=\bm{\Sigma}_{0}/s_{0}$. Furthermore, the 2nd diagonal component in (51), as shown in [10, Sect. 4] and, by using our notation in the Appendix .4, of this work is given by
@@
where $S([\mathbf{V}_{S}]_{11},\underline{\mathrm{vecs}}(\mathbf{V}_{S}))$ denotes the scale function applied to $\mathbf{V}_{S}$ reconstructed from $\mathbf{V}_{S}]_{11}$ and $\underline{\mathrm{vecs}}(\mathbf{V}_{S})$. For further reference, let us recast $\mathbf{J}[\mathbf{w}](\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0})$ as:
@@
evaluated at $\mathbf{V}_{S,0}$. As an example, we show in Appendix .5 that the term $\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}$ can be explicitly obtained for the three above-mentioned scale functions as:
@@
  • $\frac{\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{(\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m})_{1}}$ for $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$,
  • $\frac{\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{(\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m})_{1}}$ for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$.
@@
Let us now introduce the score vector for $\bm{\nu}_{0}\in\Omega$ in the model $\mathcal{P}_{\bm{\nu},g}$ in (39) as $\mathbf{s}_{\bm{\nu}_{0}}=(\mathbf{s}_{\bm{\mu}_{0}}^{T},\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{T})^{T}$. By applying the rules of the differential matrix calculus detailed in [23, Ch. 8], the Kronecker product and the vec operator [23, Ch. 3], and the stochastic representation theorem (31), we get the following expressions of the scores [22, eqs. (34) and (37)]:
@@
where, following the notation adopted in [8], we defined:  [^2] To avoid confusion, it is important to note that the function $\varphi_{0}(t)$ is linked to the equivalent function $\psi_{0}(t)$ used in [22, 24] by the constant 2-2: $\varphi_{0}(t)=-2\psi_{0}(t)$.
@@
where  [^3] Note that, in the calculations, we used the equality $E\{\mathcal{Q}\varphi_{0}(\mathcal{Q})\}=m$. In fact, from (58), (33) and (30), we have $E\{\mathcal{Q}\varphi_{0}(\mathcal{Q})\}=-2\delta_{m}^{-1}\int_{0}^{\infty}q^{m/2}dg=-2\delta_{m}^{-1}[q^{m/2}g(q)]_{0}^{\infty}+m\delta_{m}^{-1}\int_{0}^{\infty}q^{m/2-1}g(q)dq=m$.
@@
In this section, we will specialise Lemmas II.1 and II.2, introduced in Sect. II, to the study of the elliptical distributions. The semiparametric model from which we start is the “canonical” one, i.e. $\mathcal{P}_{\bm{\eta},g}$ given in (40), corresponding to (c.t.) $\mathcal{P}$ in (5), along with its parametric submodels $\mathcal{P}_{\bm{\gamma}}$ in (44) (c.t. $\mathcal{P}_{1}$ in (6)), $\mathcal{P}_{s}$ in (45) (c.t. $\mathcal{P}_{2}$ in (7)), $\mathcal{P}_{\bm{\gamma},s}$ in (46) (c.t. $\mathcal{P}_{1,2}$ in (8)) and the non-parametric model $\mathcal{P}_{g}$ in (47) (c.t. $\mathcal{P}_{3}$ in (9)). Moreover, we will investigate the link between the parametric model $\mathcal{P}_{\gamma,s,\bm{\zeta}}$, introduced in (48), with the above-mentioned models. The key result that allows us to study the impact of finite and infinite dimensional nuisance parameters on the semiparametric efficiency in $\mathcal{P}_{\bm{\eta},g}$ for the parameter vector of interest $\bm{\gamma}_{0}$ is Lemma II.2. To specialize this lemma to our context, we need to draw the following connections:
  • the finite-dimensional vector of the parameters of interest is $\bm{\gamma}_{0}=(\bm{\mu}_{0}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})^{T})^{T}$ and the related score is
@@
  • The infinite-dimensional nuisance tangent space $\mathcal{T}_{g_{0}}$ (c.t. $\mathcal{T}_{3}$) is given in Appendix .7 as:
@@
$\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})\triangleq\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\perp\mathcal{T}_{g_{0}},$ (77)
@@
It now follows from equations (80) and (81) that condition (78) is satisfied.
@@
or equivalently $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$ iff $\mathrm{vec}(\mathbf{V}_{S,0}^{-1})\in\text{Ker}\;\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})=\{a\cdot\mathrm{vec}(\mathbf{D}_{S}^{\mathbf{V}_{S,0}}),a\in\mathbb{R}\}$ where the matrix $\mathbf{D}_{S}^{\mathbf{V}_{S,0}}$ is formally defined in (157) of .6. As already noted in [10, Th. 3.1], this is a direct consequence of the Property P4 of the matrix $\mathbf{M}_{S}^{\mathbf{V}_{S,0}}$ (see appendix .6). In fact, for $S\equiv S_{d}$, we have that $\mathbf{D}_{S_{d}}^{\bm{\Sigma}_{0}}=m^{-1}|\bm{\Sigma}_{0}|^{1/m}\bm{\Sigma}_{0}^{-1}=m^{-1}\mathbf{V}_{S,0}^{-1}$. Consequently, we immediately have that $\mathrm{vec}(\mathbf{V}_{S_{d},0}^{-1})\in\text{Ker}\;\mathbf{M}_{S_{d}}^{\mathbf{V}_{S,0}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$ and then $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$. ∎
@@
$\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})\triangleq\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\perp\mathcal{T}_{\bm{\zeta}_{0}},$ (86)
@@
V Parametric and semiparametric information bounds in $\mathcal{P}_{\bm{\eta},g}$
@@
The following two subsections are organized as follows: in subsect. V-A, an explicit expression for the parametric CRB of $\bm{\gamma}_{0}$ in $\mathcal{P}_{\bm{\gamma},s}$ in the presence of the nuisance parameter $s_{0}$ is provided, while the subsect. V-B proposes the counterpart of the Propositions IV.1 and IV.2 in terms of parametric and semiparametric information bounds.
@@
To calculate (in a closed-form not containing matrix inverses) the $\mathrm{CRB}(\bm{\gamma}_{0}|s_{0})$ in (91), we will exploit the expression of the CRB on $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma})_{0}^{T})^{T}$ deduced from the inverse of the FIM $\mathbf{I}_{\bm{\nu}_{0}}$. To this end, let us consider the inverse, say $\mathbf{w}^{-1}$, of the diffeomorphism $\mathbf{w}$ given in (50). It can be explicitly obtained as:
$\begin{split}\mathbf{w}^{-1}:\;&\Omega\rightarrow\Phi\\ &\bm{\nu}=(\bm{\mu}^{T},\mathrm{vecs}(\bm{\Sigma})^{T})^{T}\mapsto\mathbf{w}^{-1}(\bm{\nu})=(\bm{\mu}^{T},S^{-1}(\bm{\Sigma})\cdot\underline{\mathrm{vecs}}(\bm{\Sigma})^{T},S(\bm{\Sigma}))^{T},\end{split}$ (94)
@@
The following proposition is proved in Appendix .8:
@@
$\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})\triangleq\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}=\alpha(g_{0})^{-1}\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\mathbf{P}_{S}^{T}(\mathbf{V}_{S,0})\mathbf{D}_{m}^{\#T}\underline{\mathbf{I}}_{m}^{T},$ (98)
@@
  • $\mathbf{I}_{m^{2}}-{\rm vec}(\mathbf{V}_{S,0}){\bf e}^{T}_{1,m^{2}}$, for $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$,
  • $\mathbf{I}_{m^{2}}-\frac{1}{m}{\rm vec}(\mathbf{V}_{S,0}){\rm vec}(\mathbf{I}_{m})^{T}$, for $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$,
  • $\mathbf{I}_{m^{2}}-\frac{1}{m}{\rm vec}(\mathbf{V}_{S,0}){\rm vec}(\mathbf{V}_{S,0}^{-1})^{T}$, for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$.